# Predictable Compression Failures Why Language Models Actually Hallucinate

### v1

@@ Abstract
Large language models perform near-Bayesian inference yet violate permutation invariance on exchangeable data. We resolve this by showing transformers minimize expected conditional description length (cross-entropy) over orderings, $\mathbb{E}_{\pi}[\ell(Y\!\mid\!\Gamma_{\pi}(X))]$, which admits a Kolmogorov-complexity interpretation up to additive constants, rather than the permutation-invariant description length $\ell(Y\!\mid\!X)$. This makes them Bayesian in expectation, not in realization. We derive (i) a Quantified Martingale Violation bound showing order-induced deviations scale as $O(\log n)$ with constants; (ii) the Expectation-level Decompression Law linking information budgets to reliability for Bernoulli predicates; and (iii) deployable planners (B2T/RoH/ISR) for answer/abstain decisions. Empirically, permutation dispersion follows $a+b\ln n$ (Qwen2-7B $b\approx 0.377$, Llama-3.1-8B $b\approx 0.147$); permutation mixtures improve ground-truth likelihood/accuracy; and randomized dose-response shows hallucinations drop by $\sim$0.13 per additional nat. A pre-specified audit with a fixed ISR=1.0 achieves near-0% hallucinations via calibrated refusal at 24% abstention. The framework turns hallucinations into predictable compression failures and enables principled information budgeting.
@@ 1 Introduction
Large language models (LLMs) exhibit a fundamental paradox: they demonstrate remarkable in-context learning capabilities consistent with Bayesian inference Xie et al. [2022], Zhang et al. [2023], Bai et al. [2023], yet systematically violate the permutation invariance that characterizes Bayesian predictive distributions on exchangeable data Falck et al. [2024]. For truly exchangeable sequences where $p(x_{1},\ldots,x_{n})=p(x_{\pi(1)},\ldots,x_{\pi(n)})$, Bayesian inference yields predictive distributions satisfying $\mathbb{E}[f(X_{n+1})|X_{1},\ldots,X_{n}]=\mathbb{E}[f(X_{n+1})|X_{\pi(1)},\ldots,X_{\pi(n)}]$ for any permutation $\pi$. The violation of this property indicates a mismatch between transformer architectures and exchangeability assumptions, yet transformers achieve near-optimal statistical performance across diverse tasks. How can models be simultaneously Bayesian and non-Bayesian?
We validate theory using observables not defined by our metrics: ground-truth likelihood/accuracy under permutation mixtures, the log-scaling of permutation-induced dispersion, and randomized dose-response of hallucination to information budgets (see Appendix B for full non-circular validation principle).
@@ 1.1 Our Contributions
We resolve this paradox through three interconnected theoretical and empirical contributions:
1. Theoretical Resolution: Conditional Complexity Minimization. We show that transformers with positional encodings minimize expected conditional description length (cross-entropy), $\mathbb{E}_{\pi}[\ell(Y|\Gamma_{\pi}(X))]$, rather than the permutation-invariant $\ell(Y|X)$. This admits a Kolmogorov-complexity interpretation up to additive constants. This fundamental distinction explains why transformers are “Bayesian in expectation, not in realization”: they achieve entropy-level compression and calibration when averaged over orderings but systematically deviate for fixed input sequences. Our Quantified Martingale Violation theorem (Theorem 1) provides an explicit $O(\log n)$ upper bound with constants for these deviations under harmonic positional decay.
2. Compression Failure Framework: From Theory to Hallucination. Building on recent connections between compression and learning Delétang et al. [2023], Kalai and Vempala [2024], we develop the Expectation-level Decompression Law (EDFL, Theorem 5) that quantifies precisely how much information $\bar{\Delta}$ is required to lift an event’s probability from prior mass $\bar{q}$ to posterior mass $p$. Unlike existing work that treats compression as binary success/failure, EDFL provides exact bounds: for fixed $p=1-\varepsilon$, achieving reliability $1-\varepsilon$ for rare events ($q\ll 1$) requires $\bar{\Delta}\geq(1-\varepsilon)\log\frac{1}{\bar{q}}+O(\bar{q})$ nats. This transforms hallucination from an unpredictable failure mode into a quantifiable consequence of information insufficiency.
3. Operational Deployment Tools. We bridge theory and practice through three metrics derived from EDFL:
  • Bits-to-Trust (B2T): Information needed for target reliability $h^{*}$
  • Risk-of-Hallucination (RoH): Achievable error given information budget
  • Information Sufficiency Ratio (ISR): Ratio determining abstention decisions
These planners align with the pre-specified ISR boundary in frontier model audits, enabling practitioners to predict hallucination risk before generation and manage it through information budgeting.
@@ 1.2 Empirical Validation
We validate our framework through three complementary experiments:
Empirical (Primary, non-circular). (i) Permutation mixtures improve ground-truth likelihood/accuracy (mixture Jensen gain). (ii) Permutation dispersion follows $a+b\ln n$ with $b\approx 0.377$ across 3,059 items. (iii) Randomized dose-response shows $\approx 0.13$ fewer hallucinations per nat.
Deployment (Secondary, calibration only). A pre-specified ISR boundary (=1.0=1.0) exhibits 96.2% boundary alignment; this is reported as a calibration check rather than as primary validation.
Ground-truth permutation analysis establishes that permutation-induced dispersion precisely follows $O(\log n)$ across models. Causal dose-response studies establish that information content, not just prompt length, determines hallucination rates. Pre-specified permutation audits on 528 held-out QA tasks prevent post-hoc selection bias.
Our work synthesizes multiple research streams. While Xie et al. [2022] established Bayesian interpretations, Falck et al. [2024] documented permutation invariance violations, and Farquhar et al. [2024] developed detection methods, no prior work explained how these phenomena coexist or provided predictive frameworks. By showing positional processing creates an inherent gap between average-case and worst-case behavior through a positional Jensen inequality, we provide the missing theoretical foundation. Rather than treating hallucinations as inevitable Kalai and Vempala [2024] or relying on post-hoc detection Farquhar et al. [2024], practitioners can now predict and prevent failures through principled information management.
@@ 2 Related Work
Bayesian interpretations, MDL, and positional encodings. A large body of work ties transformers to Bayesian inference: in-context learning as implicit Bayes Xie et al. [2022], approximate Bayesian model averaging Zhang et al. [2023], near-optimal predictive power Bai et al. [2023], and full Bayesian inference comparable to MCMC Reuter et al. [2024]. However, Falck et al. [2024] show that LLMs violate permutation invariance. The compression-learning connection has deep roots in Minimum Description Length (MDL) Grünwald [2007], with Delétang et al. [2023] establishing equivalence between language modeling and compression. Positional encodings fundamentally alter behavior. Liu et al. [2024] documented systematic “lost-in-the-middle” effects. Our $O(\log n)$ bounds with explicit constants provide finer characterization connecting to compression failure.
Hallucination detection vs prevention.  Farquhar et al. [2024] introduced semantic entropy for detection, achieving AUROC 0.790 by clustering semantically equivalent responses. Kalai and Vempala [2024] proved calibrated models must hallucinate at rates approximated by Good-Turing estimation. Multiple 2024 papers adopted “LLMs as lossy compression” perspectives Su et al. [2024], Vasilatos et al. [2024], but without formal frameworks. While these works provide detection methods or prove inevitability, our EDFL offers the first predictive theory linking compression failure to hallucination with preventive measures through information budgeting.
@@ 3.1 Setup and Notation
Let $X=(x_{1},\ldots,x_{n})$ be a sequence with sufficient statistic $T(X)$ for predicting target $Y$. Let $\pi$ denote a permutation of chunk indices $\{1,\ldots,n\}$, $\Gamma_{\pi}$ a content-preserving reordering map where $\Gamma_{\pi}(X)=(x_{\pi(1)},\ldots,x_{\pi(n)})$, and $S_{\pi}(\cdot)=\Pr(\cdot|\Gamma_{\pi}(X))$ the model’s predictive distribution given permuted input. Define $q_{\pi}(x)=\mathbb{E}[f(X_{n+1})|\Gamma_{\pi}(x)]$ as the prediction under permutation $\pi$ and $\bar{q}(x)=\mathbb{E}_{\pi}[q_{\pi}(x)]$ as the average prediction over permutations. All expectations over $\pi$ are conditional on fixed item $x$. We assume $P\ll S_{\pi}$ for all $\pi$ (absolute continuity), smoothing zero-probability tokens with $\varepsilon=10^{-9}$ in practice. We write $E_{\text{pair}}:=\mathbb{E}_{\pi,\pi^{\prime}}|q_{\pi}(x)-q_{\pi^{\prime}}(x)|$ for the expected pairwise absolute difference across independent permutations.
@@ 3.2 Quantified Martingale Violations
For the Answer/Refusal indicator $f$ and position-aware predictions $q_{\pi}(x):=\mathbb{E}[f(X_{n+1})|\Gamma_{\pi}(x)]$, the permutation residual is $R_{\pi}(x):=q_{\pi}(x)-\bar{q}(x)$ where $\bar{q}(x):=\mathbb{E}_{\pi}[q_{\pi}(x)]$.
@@
$\mathbb{E}_{\pi}\big{|}R_{\pi}(x)\big{|}\;\leq\;\frac{C_{\mathrm{tot}}(x)}{4}\times\begin{cases}\displaystyle\frac{1}{1-\alpha}\big{(}n^{1-\alpha}-1\big{)},&\alpha\in(0,1),\\[6.0pt] \displaystyle H_{n-1}\;=\;\log n-\gamma+o(1),&\alpha=1,\\[4.0pt] \displaystyle\zeta(\alpha)+o(1),&\alpha>1.\end{cases}$
@@
Suppose Assumption 2 holds with nonnegative weights $w_{i}$ summing to 11 and $\psi$ $(\alpha,C)$-regular. Then Assumption 1 holds with $\Delta_{i,t}\leq w_{i}\,C\,t^{-\alpha}$ and hence $C_{\mathrm{tot}}(x)=C$.
Theorem 2 (Quantified Martingale Violation).
@@
3.3 MDL Optimality Through Architectural Closure
@@
Under the architectural closure of Theorem 3, transformers achieve information-theoretic optimality when averaged over permutations. The risk decomposes as:
$\mathbb{E}_{X,Y,\pi}[-\log p_{\theta}(Y|\Gamma_{\pi}(X))]=H(Y|T)+\mathbb{E}_{X,\pi}\operatorname{KL}(P(\cdot|T(X))\|p_{\theta}(\cdot|\Gamma_{\pi}(X)))$
@@
$\mathfrak{J}_{\Gamma}(P,\theta):=\mathbb{E}_{\pi}[\operatorname{KL}(P\|p_{\theta}(\cdot|\Gamma_{\pi}))]-\operatorname{KL}(P\|\bar{p}_{\theta})\geq 0$
@@
$\inf_{\theta}\mathbb{E}_{(X,Y)}\mathbb{E}_{\pi}[-\log p_{\theta}(Y|\Gamma_{\pi}(X))]=H(Y|T)+o(1)$
@@ 3.4 The Expectation-level Decompression Law
Insufficient information leads to compression failure manifesting as hallucination.

### v2

# Predictable Compression Failures: Order Sensitivity and Information Budgeting for Evidence-Grounded Binary Adjudication

@@ Abstract
Transformers used for evidence-grounded question answering with binary adjudication (e.g., support/refute or yes/no) can be highly sensitive to the order in which exchangeable evidence is presented, producing dispersion across permutations and unreliable attempted answers (“hallucinations” under a Bernoulli predicate). We treat evidence order as a nuisance variable and show that next-token training minimizes expected conditional description length over orderings, which can be close to Bayes-optimal in expectation while deviating under any fixed ordering. We quantify this expectation–realization gap via a Quantified Martingale Violation (QMV) bound that predicts $O(\log n)$ growth in permutation dispersion under harmonic positional sensitivity. We then derive the Expectation-level Decompression Law (EDFL), relating expected information budget to achievable reliability for Bernoulli predicates, and use it to define Bits-to-Trust (B2T), Risk-of-Hallucination (RoH), and the Information Sufficiency Ratio (ISR) with a fixed ISR-gating rule for answer/abstain decisions under permutation mixtures. On 3,059 grounded items from a five-benchmark evidence-grounded QA suite (FEVER, HotpotQA, NQ-Open, PopQA, and Controls), we observe logarithmic dispersion and Jensen gains from uniform permutation mixtures. In a pre-specified held-out audit (528 items), an ISR=1=1 gate attains 0.0–0.7% hallucination with 20.6–27.9% abstention (95% CIs).
Evidence-grounded QA, order sensitivity, abstention
[…]
@@ 1 Introduction
Transformers with positional encodings can be strong in-context learners, yet their predictions can vary substantially under permutations of the same evidence (Falck et al., 2024; Liu et al., 2024). This is a practical problem in evidence-grounded QA: the evidence set is intended to be exchangeable (order should not matter), but the model’s probability of answering correctly (or refusing) can change when chunks are reordered. We study when such order sensitivity produces unreliable attempted answers (“hallucinations” under a binary adjudication predicate) and connect it to an analytically derived answer/abstain interface based on information budgets under permutation mixtures.
Our starting point is that next-token training minimizes cross-entropy on sequences, which corresponds to minimizing expected conditional description length over orderings. This reconciles two facts: models can be close to Bayes-optimal on average while deviating from permutation-invariant predictors for any fixed ordering. We formalize this expectation-versus-realization gap via a dispersion bound (QMV) and connect it to a deployment interface: when the evidence order induces high dispersion and information is insufficient, abstain rather than guess (Algorithm 1). The resulting ISR gate is fixed at 1 by theory and depends only on per-permutation Bernoulli probabilities (not post-hoc threshold tuning).
This paper is not a general hallucination benchmark: we focus on evidence-grounded binary adjudication (verifiable Bernoulli predicates and answer/abstain decisions) to isolate the roles of order sensitivity and information budgets.
Qwen2-7B Llama3.1 00.10.10.20.20.30.30.40.4Slope bbDispersion scaling
Qwen2-7B Llama3.1 $5\cdot 10^{-2}0$.10.1Jensen gapMixture gain (nats/token)
Figure 1: Order sensitivity and mixture gains (Experiment 1; N=3,059N=3{,}059, $m=16$ banded permutations). Left: estimated log-scaling slope $b$ in the empirical law $|R_{\pi}|\approx a+b\log n$ (95% CI as defined in Section 4.1; Table 2). Right: Jensen gap (nats/token) from uniform permutation mixtures, quantifying the cross-entropy improvement from averaging over evidence orderings.
@@ 1.1 Our Contributions
We make three contributions:
  • Quantified order sensitivity. We prove a Quantified Martingale Violation (QMV) bound controlling permutation-induced dispersion, with explicit constants and $O(\log n)$ scaling under harmonic positional sensitivity (Theorem 1 and Theorem 2).
  • EDFL as an actionable Bernoulli specialization. EDFL (Theorem 5) is a Bernoulli coarse-graining of a convexity + data-processing bound, yielding closed-form planners (B2T/RoH/ISR) and a fixed ISR=1{=}1 answer/abstain rule; we also derive a rare-event lower bound scaling like $\log(1/\bar{q})$.
  • A deployable answer/abstain rule. From EDFL we define Bits-to-Trust (B2T), Risk-of-Hallucination (RoH), and the Information Sufficiency Ratio (ISR), and give an ISR-gating procedure under permutation mixtures (Algorithm 1).
@@ 1.2 Empirical Validation
We validate the theory using observables not defined by the planners themselves (Appendix B). In particular:
  • Permutation mixtures improve log-loss and accuracy (Experiment 1), and dispersion follows $a+b\log n$ across depths (Figure 1).
  • Randomized dose-response identifies a causal effect of information on hallucination, holding prompt length fixed (Experiment 2; Figure 2).
  • A pre-specified held-out audit reports an out-of-sample operating point for ISR gating on the reliability–coverage plane without post-hoc threshold tuning (Experiment 3; Figure 3).
[…]
Figure 2: Causal dose-response of information (Experiment 2). Randomized study holding prompt length fixed at L=4L{=}4 chunks while varying support dose $d\in\{0,1,2,3\}$ (subset of Factuality Slice; Appendix H.4). We plot the net change from dose $0\to 3$ in hallucination rate against the induced decrease in the information-budget estimate ($\Delta\bar{\Delta}:=\bar{\Delta}_{d=0}-\bar{\Delta}_{d=3}\approx 3\times 0.375=1.125$ nats). The fitted slope (OLS) is 0.127 fewer hallucinations per additional nat (12.7-12.7 pp/nat); over the same shift, answer rate increases by 37.5pp and accuracy on attempts by 45.6pp.
7070727274747676787880808282848400.20.20.40.40.60.60.80.811Coverage (%)Hallucination (%)
Figure 3: Audit operating point (Experiment 3; $N=528$, $m=6$). Pre-specified held-out audit with ISR=1=1 and fixed seeds. Shaded region shows the 95% CI for coverage (1–abstention) and hallucination rate (Wilson intervals as in Section 4.1); midpoint shown for visualization.
@@ 2 Related Work
Order sensitivity and permutation invariance. Transformers can be highly sensitive to evidence order, including systematic “lost-in-the-middle” effects (Liu et al., 2024), and formal violations of permutation invariance under positional processing (Falck et al., 2024). We focus on quantifying the resulting dispersion with an explicit $O(\log n)$ bound (QMV) and connecting it to an operational deployment interface.
Bayesian/MDL views of in-context learning. Prior work connects transformers and in-context learning to Bayesian inference in various forms (Xie et al., 2022; Zhang et al., 2023) (see also (Bai et al., 2023; Reuter et al., 2024)). The compression–learning connection is classically captured by Minimum Description Length (MDL) (Grünwald, 2007), and recent work relates language modeling and compression explicitly (Delétang et al., 2023). Our contribution is to make the expectation vs. realization gap over evidence orderings explicit and measurable (QMV and Jensen gaps), rather than treating order-invariance as an implicit property of the predictor.
Hallucination control, calibration, and abstention. There is extensive work on detection and calibration signals, e.g. semantic entropy for detecting errors (Farquhar et al., 2024) and lower bounds on calibrated predictors tied to Good–Turing-like effects (Kalai and Vempala, 2024). Several recent papers frame LLM behavior through lossy-compression lenses (Su and others, 2024; Vasilatos and others, 2024). Relative to this literature, we focus on evidence-grounded Bernoulli adjudication and provide a deployment-oriented budget–reliability law (EDFL) that yields a non-tunable ISR=1=1 answer/abstain rule under permutation mixtures. Our answer/abstain interface is also related to selective classification (reject option) and risk–coverage tradeoffs (Chow, 1970; Geifman and El-Yaniv, 2017); our emphasis is on the Bernoulli coarse-graining plus permutation mixtures that yield an audited, fixed-threshold ISR boundary.
@@ 3.1 Setup and Notation
Let $x$ denote an instance consisting of $n$ evidence chunks $X=(x_{1},\ldots,x_{n})$ and a target variable $Y$. Let $\pi$ be a permutation of chunk indices $\{1,\ldots,n\}$ and let $\Gamma_{\pi}$ reorder the chunks, $\Gamma_{\pi}(X)=(x_{\pi(1)},\ldots,x_{\pi(n)})$. Let $S_{\pi}(\cdot)$ denote the model’s predictive distribution over $Y$ given $\Gamma_{\pi}(X)$. We use $P$ to denote a target distribution over $Y$ (ground truth in supervised evaluation). In deployment, we instantiate $P$ with a reference distribution $P_{\mathrm{ref}}$ induced by the application’s specification for a predicate $g$ (gold labels in supervised evaluation; deterministic verification primitives such as database lookups, unit tests, or rules; human review; or a calibrated verifier). Our guarantees are stated relative to $P_{\mathrm{ref}}$.
Importantly, we do not need to construct a reference distribution over free-form outputs. For a binary adjudication predicate $g:\mathcal{Y}\to\{0,1\}$, define the induced (pushforward) distributions on $\{0,1\}$:
$P_{\mathrm{ref},g}(z)$ $=P_{\mathrm{ref}}(g(Y)=z),$
$S_{\pi,g}(z)$ $=S_{\pi}(g(Y)=z),\qquad z\in\{0,1\}.$
By data processing (coarse-graining cannot increase KL), $\operatorname{KL}(P_{\mathrm{ref}}\|S_{\pi})\geq\operatorname{KL}(P_{\mathrm{ref},g}\|S_{\pi,g})$, so the relevant reference object for answer/abstain decisions is the Bernoulli parameter $p_{\mathrm{ref}}:=P_{\mathrm{ref},g}(1)$. For a Bernoulli predicate $g:\mathcal{Y}\to\{0,1\}$ (e.g., correctness, answer-vs-refuse, constraint satisfaction), define $q_{\pi}(x):=S_{\pi}(g{=}1)$ and $\bar{q}(x):=\mathbb{E}_{\pi}[q_{\pi}(x)]$. We assume $P_{\mathrm{ref}}\ll S_{\pi}$ for all $\pi$ (absolute continuity), smoothing zero-probability tokens with $\varepsilon=10^{-9}$ in practice. We write $E_{\text{pair}}:=\mathbb{E}_{\pi,\pi^{\prime}}|q_{\pi}(x)-q_{\pi^{\prime}}(x)|$ for the expected pairwise absolute difference across independent permutations (with $\pi,\pi^{\prime}$ i.i.d. uniform over permutations). All expectations over $\pi$ are conditional on fixed instance $x$.
@@ 3.2 Quantified Martingale Violations
For a Bernoulli predicate $g$ and position-aware predictions $q_{\pi}(x):=S_{\pi}(g{=}1)$, the permutation residual is $R_{\pi}(x):=q_{\pi}(x)-\bar{q}(x)$ where $\bar{q}(x):=\mathbb{E}_{\pi}[q_{\pi}(x)]$.
@@
$\mathbb{E}_{\pi}\big|R_{\pi}(x)\big|$ $\;\leq\;\frac{C_{\mathrm{tot}}(x)}{4}\times$
[…]
In words. QMV bounds permutation-induced dispersion by the cumulative sensitivity of the logit to adjacent rank changes. Under harmonic decay ($\alpha=1$), the bound grows like $\log n$, motivating the empirical $a+b\log n$ scaling test in Experiment 1.
@@
Suppose Assumption 2 holds with nonnegative weights $w_{i}$ summing to $1$ and $\psi$ $(\alpha,C)$-regular. Then Assumption 1 holds with $\Delta_{i,t}\leq w_{i}\,C\,t^{-\alpha}$ and hence $C_{\mathrm{tot}}(x)=C$.
Theorem 2 (Explicit log-scaling under first-order positional sensitivity).
@@
3.3 MDL Optimality Through Architectural Closure (conceptual)
How to read this subsection. The results below are representational: they formalize that finite permutation mixtures are realizable if one allows tied-weight multi-branch evaluation with an averaging head (i.e., ensembling “inside” the hypothesis class). This is a conceptual lens on expectation-versus-realization gaps and does not claim that a standard single-branch deployed transformer literally implements the multi-branch architecture or that SGD recovers the mixture without additional structure. In practice, expectation over orderings comes either from training-time exposure to randomized orderings or from inference-time permutation mixtures.
@@ 3.3 MDL Optimality Through Architectural Closure (conceptual)
Under the architectural closure of Theorem 3, the permutation-mixture family can achieve information-theoretic optimality in expectation over orderings. Write $P_{T}:=P(\cdot|T(X))$ and $p_{\theta,\pi}:=p_{\theta}(\cdot|\Gamma_{\pi}(X))$. The risk decomposes as:
$\mathbb{E}_{X,Y,\pi}\big[-\log p_{\theta,\pi}(Y)\big]$ $=H(Y|T)$
$\quad+\mathbb{E}_{X,\pi}\operatorname{KL}\!\big(P_{T}\,\|\,p_{\theta,\pi}\big)$
@@
$\mathfrak{J}_{\Gamma}(P,\theta):=\mathbb{E}_{\pi}[\operatorname{KL}(P\|p_{\theta,\pi})]-\operatorname{KL}(P\|\bar{p}_{\theta})\geq 0$
@@
$\inf_{\theta}\mathbb{E}_{(X,Y)}\mathbb{E}_{\pi}\big[-\log p_{\theta,\pi}(Y)\big]=H(Y|T)+o(1)$
@@ 3.4 The Expectation-level Decompression Law
Remark 1 (Bernoulli instantiation).
Fix a predicate $g$ and consider the induced distributions $P_{\mathrm{ref},g}$ and $S_{\pi,g}$ on $\{0,1\}$. Applying EDFL on this coarse-grained space yields the information budget
$\bar{\Delta}_{g}$ $=\mathbb{E}_{\pi}[\operatorname{KL}(P_{\mathrm{ref},g}\|S_{\pi,g})]$
$=\mathbb{E}_{\pi}\big[\operatorname{KL}(\operatorname{Ber}(p_{\mathrm{ref}})\|\operatorname{Ber}(q_{\pi}))\big],$
which depends only on Bernoulli probabilities. By data processing, $\mathbb{E}_{\pi}[\operatorname{KL}(P_{\mathrm{ref}}\|S_{\pi})]\geq\bar{\Delta}_{g}$, so working in Bernoulli space is conservative when the downstream decision is binary.
Novelty note. EDFL itself is an immediate consequence of convexity of KL and data processing (Appendix A.4). We highlight it because the Bernoulli coarse-graining yields a closed-form, deployment-facing rule (B2T/RoH/ISR) with a fixed ISR=1=1 threshold under permutation mixtures.
In words. EDFL lower-bounds the expected information budget needed to shift an event’s probability from prior mass $\bar{q}$ to posterior mass $p$. For rare events, the required budget grows like $\log(1/\bar{q})$ (Corollary 1), which motivates B2T as a minimum budget for target reliability and the ISR gate as a deployment-time rule.
@@
Insufficient information can lead to compression failure manifesting as hallucination.